OpenStudy (anonymous):

here are 18 students in Mr.Alvarez'S art class. Throughout the year every student must pair up with every other student to complete a project. How many projects will be finished? 149 153 171 125

OpenStudy (anonymous):

9? If every student has a partner, then you would divide the total amount of students by 2.

OpenStudy (anonymous):

Well one of th answers has to be 149, 153, 171, 125

hartnn (hartnn):

ok, lets assume there are only 3 students, then total no. of projects = 2+1 = 3 for 4 students, total number of projects = 3+2+1.= 6 similarly for 18 students it will be 17+16+15+.....+2+1 = ??

OpenStudy (anonymous):

51

hartnn (hartnn):

how? no.

OpenStudy (anonymous):

Oh

hartnn (hartnn):

its 17+16+15+14+13+12+11+10+9+8+7+6+5+4+3+2+1=? or u can do 17*(17+1)/2

OpenStudy (anonymous):

Oh ok I see now thanks

hartnn (hartnn):

welcome ^_^